# A First-look at Spatially-resolved Infrared Supernova remnants in M33 with JWST

### v1

@@ Abstract
We present the first spatially-resolved infrared images of supernova remnants (SNRs) in M33 with the unprecedented sensitivity and resolution of JWST. We analyze 43 SNRs in four JWST fields: two covering central and southern M33 with separate NIRCam (F335M, F444W) and MIRI (F560W, F2100W) observations, one 5 kpc-long radial strip observed with MIRI F770W, and one covering the giant HII region NGC 604 with multiple NIRCam and MIRI broad/narrowband filters. Of the 21 SNRs in the MIRI field, we found three clear detections (i.e., identical infrared and H morphologies), and six partial-detections, implying a detection fraction of 43% in these bands. In contrast, only one SNR (out of 16) is detectable in the NIRCam field. One of the SNRs, L10-080, is a potential candidate for having freshly-formed ejecta dust, based on its size and centrally-concentrated 21 m emission. Two SNRs near NGC 604 have strong evidence of molecular (H2) emission at 4.7 m, making them the farthest known SNRs with visible molecular shocks. Five SNRs have F770W observations, with the smaller younger objects showing tentative signs of emission, while the older, larger ones have voids. Multi-wavelength data indicate that the clearly-detected SNRs are also among the smallest, brightest at other wavelengths (H, radio and X-ray), have the broadest line widths (H FWHM250-350 km/s), and the densest environments. No strong correlation with star-formation histories are seen, with the clearly-detected SNRs having both high-mass (35 M⊙) and low-mass (10 M⊙) progenitors.
@@ 1 Introduction
Infrared (IR) studies of supernova remnants (SNRs) provide a unique, low-extinction window into the physics of SN explosions and the interaction of their blastwaves with the ambient medium. Mid-to-far-IR emission of SNR is generally dominated by thermal emission from collisionally-heated dust grains with a range of sizes and temperatures. Near-to-mid-IR emission in SNRs features a rich tapestry of line emission from shock-excited atomic and molecular species, prominent broad emission features from polycyclic aromatic hydrocarbons (PAHs), and in rare instances, synchrotron emission (see Dwek & Arendt, 1992; Reach et al., 2006; Tielens, 2008; Williams & Temim, 2017; Vink, 2020, for reviews).
@@
In this paper, we push the boundary of extragalactic SNR studies with JWST’s unprecedented leap in sensitivity and resolution with a new MIRI/NIRCam survey of M33, our closest actively star-forming and gas-rich dwarf spiral galaxy. JWST’s potential for new discoveries in SNR science has been demonstrated with recent observations of SN 1987A (e.g. Larsson et al., 2023; Jones et al., 2023; Arendt et al., 2023), and Cas A (Milisavljevic et al., 2024).
@@ 2.1 JWST data
As shown in Figure 2, we use archival Cycle 1 and 2 JWST datasets covering several regions of M33 in MIRI and NIRCam bands. We will henceforth refer to these as: the MIRI-field, NIRCam-field, the F770W-strip, and the NGC-604 fields. The filters used in these fields and the most common emission sources expected within the passbands are summarized in Table 1. Below we provide the basic features, spatial and filter coverage in these fields.
@@
MIRI field: The field was observed with the F560W and F2100W filters, covering a roughly 0.0950.145 deg2 (1.39 2.13 kpc2) patch of the southern spiral arm of M33. The field was constructed from a 55 dithered mosaic obtained as part of the GO 2128 program (PI: E. Rosolowsky). The observations were designed primarily with the aim of detecting embedded young massive stars and ISM dust (see Peltonen et al., 2024), although as we demonstrate later, the field contains a number of known SNRs in the region with unprecedented spatial detail in the near/mid-IR. Both F560W, centered at 5.6 m, and F2100W, centered at 21 m, are broadband filters that primarily capture emission from warm, stochastically-heated small dust grains in the interstellar medium. The full-width at half-maxima (FWHM) of the point spread function (PSF) are 0.21″(0.84 pc, for F560W) and 0.67″(2.75 pc for F2100W) at the distance of M33. A detailed view of the full MIRI field, and locations of the known SNRs (see Section 2.2), are shown in Figure 3.
Table 1: JWST filters used in our survey, and well known sources of nebular, shocked and continuum emission in these filters. Filter Pivot wavelength Wavelength Range Fields Expected major sources of emissiona (m) (m) F090W 0.90 0.791.01 3 Continuum (stars), Pa, [S III]:0.9,0.95m F187N 1.87 1.861.88 3 Pa F200W 1.99 1.762.23 3 Continuum (stars), Pa, H2 F335M 3.36 3.183.54 2,3 PAH, Continuum (stars) F444W 4.40 3.884.98 2,3 Br, [Ar VI]:4.53m, [Fe II]:4.65m, H2, Continuum (stars, dust) F470N 4.71 4.684.73 3 H2 0-0 S(9), CO F560W 5.64 5.056.17 1 Continuum (stars, dust), H2, PAH F770W 7.64 6.588.69 3,4 PAH, H2, [Ar III]:6.98m, [Ne VI]:7.65m F1130W 11.31 10.9511.67 3 PAH F1500W 15.06 13.5316.64 3 PAH F2100W 20.79 18.4823.16 1 Continuum (dust), [S III]:18.71m -
@@
NIRCam field: This F335M+F444W mosaic was obtained in parallel with the MIRI field, also as part of GO 2128 (Figure 4). The field covers an area of roughly 0.110.15 deg2 ( kpc2), at a resolution of 0.11″(0.45 pc) with F335M, and 0.14″(0.57 pc) with F444W. Both filters are sensitive to hot dust emission and spectral lines (e.g. Br- 4.05 m, H2 0-0) in the ISM, while emission in the F335M filter is sensitive to stellar continuum as well as the 3.3 m PAH emission feature arising from aromatic CH stretching modes.
@@
NGC 604: We include imaging of this active HII region as it provides the most panchromatic view of IR-visible SNRs in our survey, across multiple NIRCam and MIRI filters. These include: F090W, F187N (tracing Pa-), F200W, F335M and F444W (same as the NIRCam field), F470N (which sits on the H2 S(9) line and part of the CO fundamental band), and MIRI F770W, F1130W, F1500W, tracing various PAH emission features. The observations were obtained from the public program GO 6555 (PI: M.G. Marin). Two fields were observed: one centered on the HII region NGC 604 observed with MIRI and NIRCam, and an off field to the north observed only with NIRCam. The sky footprints of the NIRCam and MIRI observations of NGC 604 are slightly different, but on average, for both fields, they cover about (0.04∘)2 area, or roughly (0.59 kpc)2, as seen in Figure 2. The observations provide highly resolved views of the SNR and its surrounding ISM, with spatial resolution ranging from 0.03″(0.12 pc) for F090W to 0.48″(1.9 pc) for F1500W.
@@
F770W strip: We also include observations of a radial strip of M33 with the F770W filter, which we obtained as part of GO 3436 (PI: E. Rosolowsky). The strip is about 4 scale length from the center, spanning a physical area of about 5.25 kpc 0.62 kpc. The final map was constructed from a combination of 18 mosaic tiles, and has a spatial resolution (full-width half maximum) of 0.269″, or roughly 1.1 pc at the distance of M33. Similar to F335M, the F770W filter is dominated by the 7.7 m PAH feature, which arises from a combination of C-C stretching and C-H in-plane bending modes. The mosaic design was constructed with the purpose of tracing the spatial and radial variation of diffuse PAH emission on pc-scales, although we will mainly focus on the few SNRs that coincide with the strip.
Observations were reduced using the pjpipe reduction pipeline designed for the Physics at High Angular Resolution in Nearby GalaxieS (PHANGS)-JWST survey (see Lee et al., 2023; Williams et al., 2024; Peltonen et al., 2024, for details). The pjpipe package uses the official JWST pipeline (Bushouse et al., 2024) as a base, but includes some changes to the default pipeline parameters and the introduction of several custom steps that optimize the imaging of large fields filled with extended emission. Compared to the official pipeline, pjpipe notably implements different background matching approaches, a multistep destriping algorithm for NIRCam data, and a tiered set of astrometric alignment choices to produce good alignment between different JWST images of the same region. We use Gaia DR3 (Gaia Collaboration et al., 2023) sources to set the absolute astrometric alignment for all bands, and we find there are sufficient numbers of Gaia sources that have mid-IR counterparts to get good global alignment.
@@ 2.4 HST Images and Star-formation histories
For the brightest SNRs, we also make use of the much higher-resolution wide-band HST images from the PHAT Triangulum Extended Region (PHATTER) survey (Williams et al., 2021), specifically the F336W, F475W, F814W and F160W filters with a typical full-width at half-maximum of 0.07″(0.3 pc at M33)333https://hst-docs.stsci.edu/wfc3ihb/chapter-6-uvis-imaging-with-wfc3/6-6-uvis-optical-performance. Although SNRs are not broadband emitters at optical wavelengths, their strong emission lines can be picked up in the wide-band filters as we demonstrate later in Figure 5. The high angular resolution HST images of these SNRs allow a more accurate structural comparison with JWST than the lower-resolution LGGS images. The relevant spectral lines from shocks that appear within the above filter bandpasses are nicely illustrated in Figure 3 of Lin et al. (2020), based on predictions from the MAPPINGS III shock library of Allen et al. (2008). We expect the F336W filter to trace [Ne V]3426, F475W to trace [O III] 4959,5007 and H 4861, also possibly some [He II]4686 and [N I]5200, and F814W to trace lines from the Pa-series of H, and the [S III] 9069, 9532 lines. The strengths of these lines strongly depend on shock velocity and ambient density, and less strongly on factors like the magnetic field, composition and ionization states of the ambient material, and atomic physics (Allen et al., 2008; Sutherland & Dopita, 2017).
@@
We also make use of the PHATTER star-formation histories (SFHs, Lazzarini et al., 2022) to check how the visibility of the SNRs in the JWST images could be related to the recent SFH (56 Myr) in their vicinity. In this paper, we will be discussing these age ranges in terms of their equivalent stellar “mass” range, as measured by Koplitz et al. (2023) (hereafter, K23) from the PHATTER SFHs. These were interpreted by K23 as the most probable progenitor masses of the SNRs. We use the median and 68th percentile mass range of each SNR reported in Table 2 of K23, derived from conversion of the mass-weighted stellar age range in the SNR vicinity assuming single stellar evolution models. We refer the reader to K23 for the details of these measurements. Note that K23 reports two types of masses for each SNR that we use. The first set of masses, which we refer to as “grid-based”, are based on the location of the SNR in the Lazzarini et al. (2022) SFH map, where the SFHs are calculated in a grid of 100100 pc cells. The second one, which we refer to as “best-fit”, are from more localized SFHs measured by K23 using only stars within 50 pc of each SNR, after subtracting a ‘background’ SFH (from a 50pc to 1 kpc annulus centered on the SNR). The latter accounts for contamination by background stellar populations that are likely unrelated to the SNR. The lowest measurable mass, by definition, will be 7.1 M⊙ (corresponding to 56 Myr). Some SNRs have no best-fit mass above this mass-range, indicating that there is no recent star formation uniquely associated with the SNR. We will show these SNRs as having best-fit upper limit masses of 7.1 M⊙.
@@ 2.5 Methodology
Visual inspection reveals that while some SNRs can be very clearly seen in the JWST images, others are partially seen, or somewhat ambiguous against the background emission. For discussion purposes, we categorize the SNRs into one of four categories
Clearly-detected: the SNR is detected in all the filters, and the detected emission almost perfectly traces out the structure of the SNR in the H images in the regions where [S II]/H 0.4 (e.g. L10-039, Figure 5).
@@
Ambiguous (only for MIRI field): the SNR has some emission along its line of sight, but it is unclear whether the emission is from the SNR or the background or foreground ISM. This is particularly applicable to the MIRI field as the bright dust-rich ISM fills a substantial fraction of our observed area (e.g. L10-074, Figure 7).
@@
We note that this classification is somewhat subjective and user-dependent. More quantitative or citizen-science based classifications can be performed in future, but for the purposes of first presentation and discussion of images in this paper, we believe the above visual classification is adequate. Similar visual classification schemes have also been used for studying Magellanic Cloud SNRs in infrared images (e.g. Seok et al., 2013; Matsuura et al., 2022).

### v2

@@ Abstract
We present the first spatially-resolved infrared images of supernova remnants (SNRs) in M33 with the unprecedented sensitivity and resolution of JWST. We analyze 40 SNRs in four JWST fields: two covering central and southern M33 with separate NIRCam (F335M, F444W) and MIRI (F560W, F2100W) observations, one 5 kpc-long radial strip observed with MIRI F770W, and one covering the giant HII region NGC 604 with multiple NIRCam and MIRI broad/narrowband filters. Of the 21 SNRs in the MIRI (F560W+F2100W) field, we found three clear detections (i.e., identical infrared and H morphologies), and six partial-detections, implying a detection fraction of 43% in these bands. One of the SNRs in this field, L10-080, is a potential candidate for having freshly-formed ejecta dust, based on its size and centrally-concentrated 21 m emission. In contrast, only one SNR (out of 16) is detectable in the NIRCam F335M+F444W field. Two SNRs near NGC 604 have strong evidence of molecular (H2) emission at 4.7 m, making them the farthest known SNRs with visible molecular shocks. Five SNRs have F770W observations, with the smaller younger objects showing tentative signs of emission, while the older, larger ones have voids. Multi-wavelength data indicate that the clearly-detected SNRs are also among the smallest, brightest at other wavelengths (H, radio and X-ray), have the broadest line widths (H FWHM250-350 km/s), and the densest environments. No correlation between the JWST-detectability and local star-formation history of the SNRs is apparent.
@@ 1 Introduction
Infrared (IR) studies of supernova remnants (SNRs) provide a unique window into the physics of SN explosions and the interaction of their blastwaves with the ambient medium. Mid-to-far-IR (20 m) emission of SNR is generally dominated by thermal emission from collisionally-heated dust grains with a range of sizes and temperatures. Near-to-mid-IR (1-20 m) emission in SNRs features a rich tapestry of line emission from shock-excited atomic and molecular species, prominent broad emission features from polycyclic aromatic hydrocarbons (PAHs), and in rare instances, synchrotron emission (see Dwek & Arendt, 1992; Reach et al., 2006; Tielens, 2008; Williams & Temim, 2017; Vink, 2020, for reviews).
@@
In this paper, we push the boundary of extragalactic SNR studies with JWST’s unprecedented leap in sensitivity and resolution with a new MIRI/NIRCam survey of M33, our closest actively star-forming and gas-rich dwarf spiral galaxy. JWST’s potential for new discoveries in SNR science has been demonstrated with recent observations of SN 1987A (e.g. Larsson et al., 2023; Jones et al., 2023; Arendt et al., 2023), Cas A (Milisavljevic et al., 2024), and Crab Nebula (Temim et al., 2024).
@@ 2.1 JWST data
As shown in Figure 2, we use archival Cycle 1 and 2 JWST datasets covering several regions of M33 in MIRI and NIRCam bands. We will henceforth refer to these as: the MIRI-field, NIRCam-field, the North Radial Strip, and the NGC-604 fields. Basic survey details of these fields, including proposal IDs and observation dates are summarized in Table 1, and details of the filters themselves, along with the most common emission sources expected within the passbands are summarized in Table 2. Below we provide further description of the spatial and filter coverage in these fields.
@@
Southern Arm: The field was observed with the F560W and F2100W filters, covering a roughly 0.0950.145 deg2 (1.39 2.13 kpc2) patch of the southern spiral arm of M33. The field was constructed from a 55 dithered mosaic obtained as part of the GO 2128 program (PI: E. Rosolowsky). The observations were designed primarily with the aim of detecting embedded young massive stars and ISM dust (see Peltonen et al., 2024), although as we demonstrate later, the field contains a number of known SNRs in the region with unprecedented spatial detail in the near/mid-IR. Both F560W, centered at 5.6 m, and F2100W, centered at 21 m, are broadband filters that primarily capture emission from warm, stochastically-heated small dust grains in the interstellar medium. The full-width at half-maxima (FWHM) of the point spread function (PSF) are 0.21″(0.84 pc, for F560W) and 0.67″(2.75 pc for F2100W) at the distance of M33. A detailed view of the full MIRI field, and locations of the known SNRs (see Section 2.2), are shown in Figure 3.
Table 1: Fields observed with JWST in this paper. Locations of the fields are in Figure 2. From left to right, columns indicate the field name, JWST instrument, filters observed in the field, proposal ID and observation dates. Field Instrument Filters Proposal ID Obs Dates Southern Arm MIRI F560W, F2100W GO 2128 Aug 18, 2022 Center NIRCam F335M, F444W GO 2128 Aug 18, 2022 North Radial Strip MIRI F770W GO 3436 Jan 17, 2024 NGC 604 NIRCam F090W, F187N, F200W, F335M, F444W, F470N DD 6555 Jan 28, 2024 – MIRI F770W, F1130W, F1500W – Jan 26, 2024
Table 2: JWST filters used in our survey. From left to right, the columns indicate filter name, pivot wavelength (), wavelength range (), fields from Table 1 where filters were exposed and the corresponding exposure time in seconds, angular resolution in arcseconds (), physical resolution in pc () and the major sources of emission expected in these filters Filtere a a Fielde Exposure a,b c Expected major sources of emissiond (m) (m) (s) () (pc) F090W 0.9 0.791.01 3 2362 0.03 0.13 Continuum (stars), Pa, [S III]:0.9,0.95m F187N 1.87 1.861.88 3 4724 0.06 0.26 Pa F200W 1.99 1.762.23 3 2362 0.07 0.27 Continuum (stars), Pa, H2 F335M 3.36 3.183.54 2 3221 0.11 0.45 PAH, Continuum (stars) … … … 3 2362 … … … F444W 4.4 3.884.98 2 7516 0.15 0.59 Br, [Ar VI]:4.53m, [Fe II]:4.65m, H2, Continuum (stars, dust) … … … 3 2362 … … … F470N 4.71 4.684.73 3 4724 0.16 0.65 H2 0-0 S(9), CO F560W 5.64 5.056.17 1 14153 0.21 0.84 Continuum (stars, dust), H2, PAH F770W 7.64 6.588.69 3 699 0.27 1.10 PAH, H2, [Ar III]:6.98m, [Ne VI]:7.65m … … … 4 10190 … … … F1130W 11.31 10.9511.67 3 699 0.38 1.53 PAH F1500W 15.06 13.5316.64 3 699 0.49 1.99 PAH F2100W 20.79 18.4823.16 1 14153 0.67 2.75 Continuum (dust), [S III]:18.71m Note. — aBased on JWST User Documentation.
bEmpirical full-width at half-maximum.
cAssuming a distance of 840 kpc to M33 (Freedman et al., 2001).
dBased on Oliva et al. (1999); Reach et al. (2006); Neufeld et al. (2007); Allen et al. (2008); Rho et al. (2024).
e(1)–Southern Arm, (2)–Center, (3)–NGC 604, (4)–North Radial Strip
@@
Center: This F335M+F444W mosaic was obtained in parallel with the Southern Arm, also as part of GO 2128 (Figure 4). The field covers an area of roughly 0.110.15 deg2 ( kpc2), at a resolution of 0.11″(0.45 pc) with F335M, and 0.14″(0.57 pc) with F444W. Both filters are sensitive to hot dust emission and spectral lines (e.g. Br- 4.05 m, H2 0-0) in the ISM, while emission in the F335M filter is sensitive to stellar continuum as well as the 3.3 m PAH emission feature arising from aromatic CH stretching modes.
@@
NGC 604: We include imaging of this active HII region as it provides the most panchromatic view of IR-visible SNRs in our survey, across multiple NIRCam and MIRI filters. These include: F090W, F187N (tracing Pa-), F200W, F335M and F444W (same as the Center field), F470N (which sits on the H2 S(9) line and part of the CO fundamental band), and MIRI F770W, F1130W, F1500W, tracing various PAH emission features. The observations were obtained from the public program GO 6555 (PI: M.G. Marin). Two fields were observed: one centered on the HII region NGC 604 observed with MIRI and NIRCam, and an off field to the north observed only with NIRCam. The sky footprints of the NIRCam and MIRI observations of NGC 604 are slightly different, but on average, for both fields, they cover about (0.04∘)2 area, or roughly (0.59 kpc)2, as seen in Figure 2. The observations provide highly resolved views of the SNR and its surrounding ISM, with spatial resolution ranging from 0.03″(0.12 pc) for F090W to 0.48″(1.9 pc) for F1500W.
@@
North Radial Strip: We also include observations of a radial strip of M33 with the F770W filter, which we obtained as part of GO 3436 (PI: E. Rosolowsky). The strip is about 4 scale length from the center, spanning a physical area of about 5.25 kpc 0.62 kpc. The final map was constructed from a combination of 18 mosaic tiles, and has a spatial resolution (full-width half maximum) of 0.269″, or roughly 1.1 pc at the distance of M33. Similar to F335M, the F770W filter is dominated by the 7.7 m PAH feature, which arises from a combination of C-C stretching and C-H in-plane bending modes. The mosaic design was constructed with the purpose of tracing the spatial and radial variation of diffuse PAH emission on pc-scales, although we will mainly focus on the few SNRs that coincide with the strip.
Observations were reduced using the pjpipe reduction pipeline designed for the Physics at High Angular Resolution in Nearby GalaxieS (PHANGS)-JWST survey (see Lee et al., 2023; Williams et al., 2024; Peltonen et al., 2024, for details). The pjpipe package uses the official JWST pipeline (Bushouse et al., 2024) as a base, but includes some changes to the default pipeline parameters and the introduction of several custom steps that optimize the imaging of large fields filled with extended emission. Compared to the official pipeline, pjpipe notably implements different background matching approaches, a multistep destriping algorithm for NIRCam data, and a tiered set of astrometric alignment choices to produce good alignment between different JWST images of the same region. We use Gaia DR3 (Gaia Collaboration et al., 2023) sources to set the absolute astrometric alignment for all bands, and we find there are sufficient numbers of Gaia sources that have mid-IR counterparts to get good (0.05″) global alignment. We estimate this degree of alignment by measuring the relative positions of stars in the F770W compared to the Gaia-aligned HST data from Williams et al. (2021).
@@ 2.4 HST Images and Star-formation histories
For the brightest SNRs, we also make use of the much higher-resolution wide-band HST images from the PHAT Triangulum Extended Region (PHATTER) survey (Williams et al., 2021, GO 12055, 14610; PI: J. Dalcanton), specifically the WFC3 UVIS F336W, ACS WFC F475W and F814W filters, and WFC3 IR F160W filters with a typical full-width at half-maximum of 0.07″(0.3 pc at M33)333https://hst-docs.stsci.edu/wfc3ihb/chapter-6-uvis-imaging-with-wfc3/6-6-uvis-optical-performance. Although SNRs are not broadband emitters at optical wavelengths, their strong emission lines can be picked up in the wide-band filters as we demonstrate later in Figure 5. The high angular resolution HST images of these SNRs allow a more accurate structural comparison with JWST than the lower-resolution LGGS images. The relevant spectral lines from shocks that appear within the above filter bandpasses are nicely illustrated in Figure 3 of Lin et al. (2020), based on predictions from the MAPPINGS III shock library of Allen et al. (2008). We expect the F336W filter to trace [Ne V]3426, F475W to trace [O III] 4959,5007 and H 4861, also possibly some [He II]4686 and [N I]5200, and F814W to trace lines from the Pa-series of H, and the [S III] 9069, 9532 lines. The strengths of these lines strongly depend on shock velocity and ambient density, and less strongly on factors like the magnetic field, composition and ionization states of the ambient material, and atomic physics (Allen et al., 2008; Sutherland & Dopita, 2017).
@@
We also make use of the PHATTER star-formation histories (SFHs, Lazzarini et al., 2022) to check how the visibility of the SNRs in the JWST images could be related to the recent SFH (56 Myr) in their vicinity. In this paper, we will be discussing these age ranges in terms of their equivalent stellar “mass” range, as measured by Koplitz et al. (2023) (hereafter, K23) from the PHATTER SFHs. These were interpreted by K23 as the most probable progenitor masses of the SNRs. We use the median and 68th percentile progenitor mass range of each SNR reported in Table 2 of K23, derived from conversion of the mass-weighted stellar age range in the SNR vicinity assuming single stellar evolution models. We refer the reader to K23 for the details of these measurements. Note that K23 reports two types of progenitor masses for each SNR. The first set of masses, which we refer to as “grid-based”, are based on the location of the SNR in the Lazzarini et al. (2022) SFH map, where the SFHs are calculated in a grid of 100100 pc cells. The second one, which we refer to as “best-fit”, are from more localized SFHs measured by K23 using only stars within 50 pc of each SNR, after subtracting a ‘background’ SFH (from a 50pc to 1 kpc annulus centered on the SNR). The latter accounts for contamination by background stellar populations that are likely unrelated to the SNR. The lowest measurable progenitor mass, by definition, will be 7.1 M⊙ (corresponding to 56 Myr). Some SNRs have no best-fit progenitor mass above this mass-range, indicating that there is no recent star formation uniquely associated with the SNR. We will show these SNRs as having best-fit upper limit masses of 7.1 M⊙.
@@ 2.5 Methodology
Visual inspection reveals that while some SNRs can be very clearly seen in the JWST images, others are partially seen, or somewhat ambiguous against the background emission. For discussion purposes, we categorize the SNRs into one of four categories (primarily for discussion of the Southern Arm and Center field SNRs),
Clearly-detected: the SNR is detected in all the filters in the field, and the detected emission almost perfectly traces out the structure of the SNR in the H images in the regions where [S II]/H 0.4 (e.g. L10-039, Figure 5).
@@
Ambiguous (only for Southern Arm field): the SNR has some emission along its line of sight, but it is unclear whether the emission is from the SNR or the background or foreground ISM. This is particularly applicable to the Southern Arm field as the bright dust-rich ISM fills a substantial fraction of our observed area (e.g. L10-074, Figure 7).
@@
We note that this classification is somewhat subjective and user-dependent, and as shown later, dependent on the infrared wavelength range as well. More quantitative or citizen-science based classifications can be performed in future, but for the purposes of first presentation and discussion of images in this paper, we believe the above visual classification is adequate. Similar visual classification schemes have also been used for studying Magellanic Cloud SNRs in infrared images (e.g. Seok et al., 2013; Matsuura et al., 2022).